OpenStudy (anonymous):

How do you solve an equation where the function is on both sides? \[2f(x)-3f(\frac{1}{x}) = x^2\] Find f(2) I'm stucked at \[\frac{2f(\frac{1}{x})}{2} + \frac{x^2}{2}\]

OpenStudy (anonymous):

you are trying to find a formula for f correct?

OpenStudy (anonymous):

Yes, I meant to write. f(x) = \frac{2f(\frac{1}{x})}{2} + \frac{x^2}{2}

OpenStudy (anonymous):

\[f(x) = \frac{2f(\frac{1}{x})}{2} + \frac{x^2}{2}\]

OpenStudy (anonymous):

gimmick is to replace x by 1/x again and get two equations

OpenStudy (anonymous):

How would I calculate f(2)? \[f(2) = \frac{2f(\frac{1}{2})}{2} + \frac{2^2}{2}\] And \[f(0.5) = \frac{2f(\frac{1}{0.5})}{2} + \frac{0.5^2}{2}\]

OpenStudy (anonymous):

oh wait this is easier than i thought. i thought you were trying to find an expression for \[f(x)\] but you really just want \[f(2)\] right? so replace x by 2 and get \[2f(2)-3f(\frac{1}{2})=4\]

OpenStudy (anonymous):

now replace x by 1/2 and get \[2f(\frac{1}{2})-3f(2)=\frac{1}{4}\]

OpenStudy (anonymous):

and you want to solve for \[f(2)\] so maybe put \[y=f(2), z = f(\frac{1}{2})\] and solve the system of equations \[2y-3z=4\] \[2z-3y=\frac{1}{4}\] for y and z (you will get them both, even though you only need y) good from there?

OpenStudy (anonymous):

Perfect thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

myininaya is impressed too, i can tell!

myininaya (myininaya):

to tell you the truth i don't get it

OpenStudy (anonymous):

really? you are given the identity \[2f(x)-3f(\frac{1}{x}) = x^2\] and in fact you can use this to find an expression for \[f(x)\] but in this case the question was find \[f(2)\]

OpenStudy (anonymous):

so replace x by 2, and the replace x by 1/2, and you get two equations containing f(2) solve the system and you get it

myininaya (myininaya):

ok ok i see it now lol

OpenStudy (anonymous):

helps if you have seen a similar problem before. actually i think i answered this here a month or two ago, or at least a the same type of problem

myininaya (myininaya):

\[2f(2)-3f (\frac{1}{2})=2^2\] \[2f(\frac{1}{2})-3f(\frac{1}{\frac{1}{2}})=x^2\]

myininaya (myininaya):

oops i forgot to replace x with 1/2

myininaya (myininaya):

i have seen system of linear equations but i haven't seen one like this

myininaya (myininaya):

gj satellite

OpenStudy (anonymous):

try it with x and 1/x instead of 2 and 1/2 and see if you can find the expression for f

myininaya (myininaya):

me?

OpenStudy (anonymous):

why not? i am doing it with scratch paper. bone up on algebra is all it is

OpenStudy (anonymous):

\[2f(x)-3f(\frac{1}{x})=x^2\] \[-3f(x)+2f(\frac{1}{x})=\frac{1}{x^2}\] \[4f(x)-6f(\frac{1}{x})=2x^2\] \[-9f(x)+6f(\frac{1}{x})=\frac{3}{x^2}\] add to get \[-5f(x)=2x^2+\frac{3}{x^2}=\frac{4x^2+3}{x^2}\] solve to get \[f(x)=-\frac{4x^2+3}{5x^2}\]

OpenStudy (anonymous):

what are these "user study groups"? you know?

myininaya (myininaya):

\[2 f(x)-3 f(\frac{1}{x})=x^2\] \[2 f(\frac{1}{x})-3f(\frac{1}{\frac{1}{x}})=(\frac{1}{x})^2 =>2 f(\frac{1}{x})-3f(x)=\frac{1}{x^2} \] let f(x)=z and let f(1/x)=t equation 1 is: \[2z-3t=x^2\] equation 2 is: \[2t-3z=\frac{1}{x^2}\] solve equation 1 for z: \[z=\frac{x^2+3t}{2}\] now plug this z into equation 2: \[2t-3(\frac{x^2+3t}{2})=\frac{1}{x^2}\] now we solve this for t: \[2t-\frac{3}{2}(x^2+3t)=\frac{1}{x^2} => (2-\frac{9}{2})t-\frac{3}{2}x^2=\frac{1}{x^2}\] \[=>\frac{-5}{2}t=\frac{1}{x^2}+\frac{3}{2}x^2\]=>\[t=-\frac{2}{5x^2}-\frac{3}{5}x^2\] okay so we know t now we will take this t and plug into the expression we have solve for z: \[z=\frac{x^2+3t}{2}\] \[=\frac{x^2+3 \cdot (-\frac{2}{5x^2}-\frac{3}{5} x^2)}{2}\] \[=\frac{x^2}{2}+\frac{3}{2}(-\frac{2}{5x^2})+\frac{3}{2}(\frac{-3}{5}x^2)\] \[=\frac{x^2}{2}-\frac{3}{5x^2}-\frac{9}{10}x^2\] \[=-\frac{3}{5x^2}-\frac{4}{10}x^2=-\frac{3}{5x^2}-\frac{2}{5}x^2\] so recall we let f(x)=z and let f(1/x)=t this means we have \[f(x)=-\frac{3}{5x^2}-\frac{2}{5}x^2\] and \[f(\frac{1}{x})=-\frac{2}{5x^2}-\frac{3}{5}x^2\]

myininaya (myininaya):

satellite i don't you combine fractions right

myininaya (myininaya):

oops i missed the word think

myininaya (myininaya):

i went the long way

OpenStudy (anonymous):

i think i am write. i think

OpenStudy (anonymous):

and one reason i think it is that i did it for x = 2, got an answer, and then found \[f(2)\] and got the same number

OpenStudy (anonymous):

oh but i totally wrote the wrong thing! you are right

OpenStudy (anonymous):

i wrote \[f(x)=-\frac{4x^2+3}{5x^2}\] when it should have been \[f(x)=-\frac{2x^4+3}{5x^2}\]

OpenStudy (anonymous):

which is what you have exactly. i claim "typo" as usual